OpenStudy (anonymous):

medal and fan for best answer :):):):):) in the function f(x)=4(x^2-6x+ ) +20 what number belongs in the blank to complete the square?

OpenStudy (anonymous):

(-6/2)^2=9

OpenStudy (accessdenied):

The already-completed square will look like this: \(4 (x - a)^2 + 20 \) Let's just look at the smaller picture though. \( (x - a)^2 \) If we expand that form, we get: \(x^2 - 2a x + a^2 = x^ 2 - 6x + \_\_ \) Notice the -2a x and the -6x are corresponding terms. So -2a = -6. Then the blank corresponds with our a^ 2 part. In total, our general process is dividing the x-coefficient by 2 (a = -6/-2) and then squaring the result. (a^ 2)

OpenStudy (accessdenied):

** side note: do not be confused by the division by -2 vs. 2, we simply square the result to remove the sign entirely.

OpenStudy (anonymous):

is the answer 9?

OpenStudy (accessdenied):

Yes. :)